the position of a particle moving along the x axis is given by x =t^2-2t^3-6,where x is in meters and t in second. What is the position of the particle when it achieves its maximum speed in the positive x direction?
1
Expert's answer
2021-07-27T16:19:02-0400

Let us determine the moment of the maximum positive speed.

"v_x = x' = (t^2 - 2t^3-6)' = 2t - 6t^2."

The maximum velocity corresponds to "v_x' = 0," and "v_x' = 2 - 12t." Therefore, "v_x' = 0" when "t = \dfrac16." The velocity is "v_x = 2\cdot\dfrac16 - 6\cdot\left(\dfrac16\right)^2 = \dfrac16" and the position is "x= \left(\dfrac16\right)^2 - 2\left(\dfrac16\right)^3 - 6 \approx -5.98."